# APTx Neuron: A Unified Trainable Neuron Architecture Integrating Activation and Computation

### v2

@@ Abstract
We propose the APTx Neuron, a novel, unified neural computation unit that integrates non-linear activation and linear transformation into a single trainable expression. The APTx Neuron is derived from the APTx activation function, thereby eliminating the need for separate activation layers and making the architecture both computationally efficient and elegant. The proposed neuron follows the functional form $y=\sum_{i=1}^{n}((\alpha_{i}+\tanh(\beta_{i}x_{i}))\cdot\gamma_{i}x_{i})+\delta$, where all parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ are trainable. We validate our APTx Neuron-based architecture on the MNIST dataset, achieving up to 96.69% test accuracy in just 20 epochs using approximately 332K trainable parameters. The results highlight the superior expressiveness and computational efficiency of the APTx Neuron compared to traditional neurons, pointing toward a new paradigm in unified neuron design and the architectures built upon it.
Keywords:
APTx Neuron APTx Activation Function Activation Function Unified Neuron Deep Learning MNIST
@@ 1 Introduction
Here, $\alpha$, $\beta$, and $\gamma$ are learnable parameters. This formulation enables adaptation of the dynamic shape during training for the APTx activation function.
Furthermore, the APTx activation function can generate the SWISH($x,\rho$) activation function at parameters $\alpha=1,\beta=\rho/2,\text{and }\gamma=1/2$. Similarly, we can use the values $\alpha=1,\beta=1/2,\text{and }\gamma=1/2$ for the negative part, and $\alpha=1,\beta=1,\text{and }\gamma=1/2$ for the positive part if we want to closely approximate the MISH activation function.
@@ 3 APTx Neuron
The APTx Neuron is a novel computational unit that unifies linear transformation and non-linear activation into a single, expressive formulation. Inspired by the parametric APTx activation function, this neuron architecture removes the strict separation between computation and activation, allowing both to be learned as a cohesive entity. It is designed to enhance representational flexibility while reducing architectural redundancy.
@@ 3.2 Relationship to Traditional Neurons and Activations
    which resembles the form of a conventional linear neuron with learnable weights and bias.
@@ 3.4 Parameter Overhead and Efficiency
Each APTx Neuron introduces $13n+$ 1 trainable parameters for an input dimension $n$, compared to $n+1$ in a standard neuron. However, due to their richer expressiveness, fewer APTx Neurons or layers may be required to achieve comparable or better performance. This often leads to a favorable trade-off between parameter count and model accuracy.
@@
  • Hybrid schemes: Certain parameters may be fixed. For instance, setting $\alpha_{i}=1$ (non-trainable), while keeping $\beta_{i}$, $\gamma_{i}$, and $\delta$ trainable. For a neuron with input dimension $n$, this reduces the total number of trainable parameters from $13n+$ 1 to $12n+$ 1.
@@ 4.1 Neural Network Design
The source code for the APTx Neuron-based architecture and MNIST [3] experiment, implemented in Python [12] using the PyTorch library [11], is available at our GitHub repository [7].
@@ 6 Conclusion
This work introduced the APTx Neuron, a unified, fully trainable neural unit that integrates linear transformation and non-linear activation into a single expression, extending the APTx activation function. By learning per-input parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ for each input $x_{i}$ within a neuron, the APTx Neuron removes the need for separate activation layers and enables fine-grained input transformation. APTx Neuron generalizes traditional neurons and activations, offering greater representational power. Our MNIST experiments show that a fully connected APTx Neuron-based feedforward neural network achieves $96.69\%$ test accuracy in 20 epochs with approximately 332K trainable parameters, demonstrating rapid convergence and high efficiency. This design lays the groundwork for extending APTx Neurons to CNNs and transformers, paving the way for more compact and adaptive deep learning architectures.
@@ 6.0.1 Competing Interest
The source code and related experimental data used in this study are publicly available at: https://github.com/mr-ravin/APTxNeuron.

### v4

@@ Abstract
We propose the APTx Neuron, a novel, unified neural computation unit that integrates non-linear activation and linear transformation into a single trainable expression. The APTx Neuron is derived from the APTx activation function, thereby eliminating the need for separate activation layers and making the architecture both computationally efficient and elegant. The proposed neuron follows the functional form $y=\sum_{i=1}^{n}((\alpha_{i}+\tanh(\beta_{i}x_{i}))\cdot\gamma_{i}x_{i})+\delta$, where all parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ are trainable. We validate our APTx Neuron-based architecture on the MNIST dataset, achieving up to $96.69\%$ test accuracy within 11 epochs using approximately 332K trainable parameters. The results highlight the superior expressiveness and computational efficiency of the APTx Neuron compared to traditional neurons, pointing toward a new paradigm in unified neuron design and the architectures built upon it. Source code is available at https://github.com/mr-ravin/aptx_neuron.
@@ 1 Introduction
In Equation 1, $\alpha$, $\beta$, and $\gamma$ are learnable parameters. This formulation enables adaptation of the dynamic shape during training for the APTx activation function.
Notably, the APTx activation function subsumes several well-known activations as special cases or close approximations. For example, it exactly recovers the SWISH($x,\rho$) function when $\alpha=1$, $\beta=\rho/2$, and $\gamma=1/2$. Similarly, it can closely approximate the MISH function by setting $\alpha=1$, $\beta=1/2$, and $\gamma=1/2$ for the negative domain, and $\alpha=1$, $\beta=1$, and $\gamma=1/2$ for the positive domain.
The ReLU activation function can be approximated using the APTx activation function by fixing $\alpha=1$ and $\gamma=1/2$, with the approximation improving as $\beta$ increases and converging to ReLU in the limit $\beta\to\infty$. In practice, setting $\alpha=1$, $\beta\approx 10^{6}$, and $\gamma=1/2$ already produces a close approximation of ReLU.
@@ 3 APTx Neuron
The APTx Neuron is a novel computational unit that unifies linear transformation and non-linear activation into a single, expressive formulation. Inspired by the parametric APTx activation function, this neuron architecture removes the strict separation between computation and activation, allowing both to be learned as a cohesive entity. It is designed to enhance representational flexibility while reducing architectural redundancy. A key theoretical property of the APTx Neuron is that it naturally retains the universal approximation capability [4].
@@ 3.2 Relationship to Traditional Neurons and Activations
    The form presented in Equation 6 is structurally identical to a conventional linear neuron with learnable weights and bias. Consequently, the universal approximation property [4] is inherently preserved in the APTx Neuron.
@@ 3.4 Parameter Overhead and Efficiency
Each APTx Neuron introduces $3n+1$ trainable parameters for an input dimension $n$, compared to $n+1$ in a standard neuron. However, due to their richer expressiveness, fewer APTx Neurons or layers may be required to achieve comparable or better performance. This often leads to a favorable trade-off between parameter count and model accuracy.
@@
  • Hybrid schemes: Certain parameters may be fixed. For instance, setting $\alpha_{i}=1$ (non-trainable), while keeping $\beta_{i}$, $\gamma_{i}$, and $\delta$ trainable. For a neuron with input dimension $n$, this reduces the total number of trainable parameters from $3n+1$ to $2n+1$.
@@ 4.1 Neural Network Design
The source code for the APTx Neuron-based architecture and the MNIST [3] experiment, implemented in Python [14] using the PyTorch library [13], is available at our GitHub repository [9]. In addition, a pip-installable Python package implementing the APTx Neuron is available on both GitHub [7] and PyPI [8]. It can be installed using the following command: pip install aptx_neuron.
@@ 6 Conclusion
This work introduced the APTx Neuron, a unified, fully trainable neural unit that integrates linear transformation and non-linear activation into a single expression, extending the APTx activation function. By learning per-input parameters $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ for each input $x_{i}$, and a shared bias term $\delta$ within a neuron, the APTx Neuron removes the need for separate activation layers and enables fine-grained input transformation. The APTx Neuron generalizes traditional neurons and activations, offering greater representational power. Our experiments show that a fully connected APTx Neuron-based feedforward neural network achieves $96.69\%$ test accuracy on the MNIST dataset within 11 epochs using approximately 332K trainable parameters, demonstrating rapid convergence and high efficiency. This design lays the groundwork for extending APTx Neurons to CNNs and transformers, paving the way for more compact and adaptive deep learning architectures.
@@ 6.0.1 Competing Interest
The source code and experimental results for this study are available at: https://github.com/mr-ravin/APTxNeuron. In addition, a pip-installable Python package implementing the APTx Neuron is available on both GitHub and PyPI:
The package can be installed using the following command: pip install aptx-neuron